Need to end up with a formula set to a. I have tried to many things, but it's never right.
-1/2a(t^2) = a*t*z - 1/2g(z)

I've always struggled with being able to solve these when the piece needed is embedded in both sides :(

Answer by stanbon(75887)

(-1/2)a(t^2) = a*t*z - (1/2)g(z)
----------------------
(-1/2)a(t^2) - a*t*z = - (1/2)g(z)
----------------------
a[(-1/2)t^2 -tz] = -(1/2)g(z)
---
a = [(1/2)g(z)]/[(1/2)t^2 + tz]
==================================
Cheers,
Stan H.